OpenStudy (anonymous):

I really need help please! An element has three naturally occurring isotopes. Use the information below to calculate the weighted average atomic mass of the element, showing both the setup and the final answer for the calculation. Isotope Atomic Mass Percent Abundance X 1.01 u 99.984% Y 2.01 u 0.014% Z 3.02 u 0.002%

OpenStudy (anonymous):

Idk how to insert this into Wolfram Alpha

OpenStudy (bigdaddyharsha):

average atomic mass of the element= 1.01(0.99984) + 2.01(0.00014) + 3.02(0.00002) = 1.01 u You could make this easier on yourself and do it this way: average atomic mass of the element= 1.01(99.984/100) + 2.01(0.014/100) + 3.02(0.002/100) = 1.01 u

OpenStudy (anonymous):

Thanks youu